OpenStudy (mathmath333):

find the remainder of \(\large\tt \color{black}{50^{51^{52}}}\) divided by 11

OpenStudy (ikram002p):

hmm i wish there is new NT problems

ganeshie8 (ganeshie8):

I want to avoid using NT stuff as much as possible here because this question is for entrance tests : First, notice that 51^52 leaves a remainder of 1 when divided by 10

OpenStudy (mathmath333):

yes

ganeshie8 (ganeshie8):

you can figure that out by using binomial theorem : 51^52 = (50+1)^52 = 10M + 1

OpenStudy (mathmath333):

yes

ganeshie8 (ganeshie8):

50^51^52 = 50^(10M+1) agree ?

OpenStudy (mathmath333):

yes

ganeshie8 (ganeshie8):

if you heard of Fermat little theorem before : 50^10 leaves a remainder of 1 when `divided by 11` 50^51^52 = 50^(10M+1) = 50*50^(10M) = 50 * 1 = 50 = 6 so the remainder is 6

ganeshie8 (ganeshie8):

we need to apply binomial theorem again if you don't want to use Fermat little theorem

OpenStudy (mathmath333):

fermat rule says 50^(11n) mod 11=1 ?

OpenStudy (mathmath333):

a^kn mod k=1

OpenStudy (mathmath333):

it should be a^(k-1) mod k congruent to 1

OpenStudy (mathmath333):

ok thanks very much

OpenStudy (ikram002p):

so u got it ?

OpenStudy (mathmath333):

\(\large\tt \color{black}{a^{p-1}\bmod{p}\equiv 1\bmod{p}}\)

OpenStudy (mathmath333):

right?

OpenStudy (mathmath333):

i think i got it